The length of a triangle is three times the width. The perimeter is 44 cm. What are the dimensions of the rectangle?
Thanks for all your help :)

Answer by checkley79(3341)

L=3W
P=2L+2W
44=2*3W+2W
44=6W+2W
44=8W
W=44/8
W=5.5 cm. ANS FOR THE WIDTH.
L=3*5.5=16.5 cm. ANS FOR THE LENGTH.
PROOF:
44=2*5.5+2*16.5
44=11+33
44=44